# RLAIF-V: Aligning MLLMs through Open-Source AI Feedback for Super GPT-4V Trustworthiness

### v1

@@ Abstract
Learning from feedback reduces the hallucination of multimodal large language models (MLLMs) by aligning them with human preferences. While traditional methods rely on labor-intensive and time-consuming manual labeling, recent approaches employing models as automatic labelers have shown promising results without human intervention. However, these methods heavily rely on costly proprietary models like GPT-4V, resulting in scalability issues. Moreover, this paradigm essentially distills the proprietary models to provide a temporary solution to quickly bridge the performance gap. As this gap continues to shrink, the community is soon facing the essential challenge of aligning MLLMs using labeler models of comparable capability. In this work, we introduce RLAIF-V, a novel framework that aligns MLLMs in a fully open-source paradigm for super GPT-4V trustworthiness. RLAIF-V maximally exploits the open-source feedback from two perspectives, including high-quality feedback data and online feedback learning algorithm. In order to better expose the difference in trustworthiness, we propose a novel deconfounded candidate response generation strategy. To enhance the accuracy of pairwise feedback data from open-source MLLMs, we employ a divide-and-conquer approach that breaks down the response into atomic claims, simplifying the task to obtain more reliable results. For the feedback learning algorithm, RLAIF-V mitigates the distribution shift problem of vanilla direct preference optimization in an online learning fashion, improving both the learning performance and efficiency. Extensive experiments on seven benchmarks in both automatic and human evaluation show that RLAIF-V substantially enhances the trustworthiness of models without sacrificing performance on other tasks. Using a 34B model as labeler, RLAIF-V 7B model reduces object hallucination by 82.9% and overall hallucination by 42.1%, outperforming the labeler model. Remarkably, RLAIF-V also reveals the self-alignment potential of open-source MLLMs, where a 12B model can learn from the feedback of itself to achieve less than 29.5% overall hallucination rate, surpassing GPT-4V (45.9%) by a large margin. The results shed light on a promising route to enhance the efficacy of leading-edge MLLMs.
@@ 1 Introduction
Recent advances in multimodal large language models (MLLMs) mark a significant milestone in AI research [4, 10, 31, 32, 33, 57]. These models are trained on large-scale multimodal corpora and possess profound world knowledge, showing remarkable capabilities in tackling diverse multimodal tasks [25, 36, 42]. However, it has been commonly noticed that MLLMs are prone to confidently generating incorrect content that deviates from human preferences [49, 70, 58, 18]. In order to align MLLMs with human preferences, reinforcement learning from human feedback (RLHF) has been widely used and demonstrates substantial results [49, 58]. However, RLHF depends heavily on labor-intensive human annotations, and consequently is hard to cover the widespread misalignment between model and human preferences. Recently, reinforcement learning from AI feedback (RLAIF), which uses the preference collected from labeler models as a proxy of human preference, has shown promising potential as an alternative to RLHF [22].
Existing RLAIF methods, demonstrated at the top of Figure 1 (a), rely on ultra-powerful proprietary models to distill feedback from [26, 67, 71, 69]. Such approaches present scalability issues due to dependence on expensive API access. More critically, this paradigm essentially distills the capability of proprietary models to provide a temporary solution to bridge the performance gap between open-source models and proprietary models. As this gap continues to shrink, the community is soon facing the challenge of aligning MLLMs using labeler models of comparable capability, demonstrated at the bottom of Figure 1 (a). However, simply changing the labeler model from a proprietary model to an open-source model leads to unsatisfactory feedback regarding both data efficiency and pairwise data accuracy [6]. Moreover, even supplied with high-quality feedback, existing training methods are prone to saturate and cannot fully utilize the data [49, 1] due to the distribution shift problem, where the preference data is static during training while model output distribution is constantly shifting.
@@
To address the challenges, we propose the RLAIF-V framework, which aligns MLLMs using open-source feedback, achieving more trustworthy behavior compared with the labeler and even GPT-4V [42] (Figure 1 (b)). As shown in Figure 2, RLAIF-V leverages two key innovations to achieve this objective: (1) At the feedback quality level, we propose a novel deconfounded candidate response generation strategy for better data efficiency and a divide-and-conquer approach for higher pairwise data accuracy. The deconfounded strategy accurately exposes the genuine trustworthiness difference within response pairs by generating candidate responses from multiple sampling decoding trials under the same condition (i.e., same inputs and decoding parameters). Consequently, confounding factors such as the text style are eliminated, and the feedback is ensured to focus on the substantial content of responses. The divide-and-conquer approach decomposes the difficult response-evaluation task into simpler claim-evaluation. We divide the full response into claims to acquire scores, and then combine them to get the final assessment. It substantially simplifies the task and thus obtains more accurate feedback from open-source MLLMs. (2) For the training method, we devise an iterative alignment framework to approximate online training and mitigate the distribution shift problem of widely used direct preference optimization (DPO) [44]. Specifically, the feedback is periodically refreshed based on the latest model weights to reduce the distribution deviation. At each iteration, we update the model using the fresh feedback with direct preference optimization.
@@
Comprehensive experiments on seven benchmarks show that RLAIF-V can substantially enhance the trustworthiness of models without any human or proprietary model intervention. Using feedback from LLaVA-NeXT 34B [32] to train a LLaVA 1.5 7B [31], we significantly reduce the object hallucination rate of the base model on Object HalBench [58] by 81.0%, surpassing the labeler model by a large margin. Pushing the limit to an extreme scenario where no stronger models are available, we align OmniLMM [43] with the model itself as the labeler. Experimental results show that RLAIF-V can reduce object hallucination by 71.0% relative points in Object HalBench and reduce overall hallucination by 41.8% relative points in MHumanEval, surpassing GPT-4V by a large margin and revealing the self-alignment potential of open-source MLLMs.
@@
Figure 2: Overview of the RLAIF-V framework. (1) Given the input image and prompt, multiple candidate responses are generated with the deconfounded strategy. (2) Each response is split into atomic claims and assigned with trustworthiness scores separately by an open-source MLLM. (3) During preference learning, the model is aligned under an iterative alignment framework which periodically updates the feedback.
@@
The contribution of this work can be summarized as fourfold: (1) We present RLAIF-V, a novel framework that aligns MLLMs with open-source feedback. (2) We propose a novel deconfounded response generation strategy and a divide-and-conquer response evaluation approach to improve the feedback data quality. (3) We employ an iterative alignment method to train the model, improving both the learning efficiency and performance. (4) We conduct comprehensive experiments to demonstrate the effectiveness of the proposed framework, achieving state-of-the-art performance in trustworthiness among both open-source and proprietary MLLMs. All codes, data, and model weights are released.
@@ 2 RLAIF-V
In this section, we first elaborate on how to collect high-quality AI feedback from open-source MLLMs by introducing the response collection and evaluation process. Then we introduce the iterative alignment approach to mitigate the distribution shift problem and thus improve the scalability. See Figure 2 for an overview of the RLAIF-V framework.
@@ 2.2 Response Evaluation
Combine. After obtaining the quality assessment of each claim, we finally combine them into the score of the whole response. For each response, we denote the number of claims having $p_{no}>p_{yes}$ as $n_{rej}$, measuring how many incorrect claims are recognized by the labeler model. We use $-n_{rej}$ as the final score $S$ of the response, where a higher score indicates less incorrectness of the content. Given the score of each response, we can now construct a preference dataset for training. For each instruction $x$, we keep all response pairs $(y,y^{\prime})$ such that $S>S^{\prime}$ and choose the higher score response $y$ as the preferred response. To save the training cost, we randomly sample at most 2 pairs for each instruction and we find such filtering process only causes minor performance drop. We also investigate the effectiveness of more combining methods (see Appendix C.1).
2.3 Iterative Alignment
@@
DPO is widely used to align MLLMs with human preference. However, naive DPO faces the distribution shift problem, i.e., the preference data is static during training process while model output distribution is constantly shifting [14]. As a result, the data distribution might deviate from the expected feedback distribution and cause sub-optimal alignment results.
Mitigating the distribution shift problem requires updating the feedback data constantly during the training process. However, collecting new feedback data at each optimization step is costly and makes the training unstable. We tackle this problem by conducting the alignment process including both data collection and training in an iterative manner. The iterative alignment is presented in Algorithm 1.
@@
Specifically, we select $N$ multimodal instructions at the beginning of each iteration and leverage the deconfounded strategy to generate $n$ candidate responses for each instruction with the latest instruction model $M_{i}$. We assign each response with a trustworthiness score through the divide-and-conquer approach using the labeler model $L$ and construct comparison pairs $D_{i}$ for training. Then we train the $M_{i}$ with direct preference optimization on $D_{i}$ to get $M_{i+1}$, which is used as the instruction model of next iteration. In this way, the feedback distribution can be updated in an iterative manner, resulting in better learning efficiency.
Algorithm 1 Iterative alignment of RLAIF-V
Initial instruction model $M_{1}$, labeler model $L$, instruction sets $X=\{X_{1},X_{2},\dots,X_{k}\}$, each $X_{i}$ contains $N$ different multimodal instructions
$M_{k}$
for $i=1$ to $k$ do
     Select the instruction set $X_{i}\in X$ of current iteration
     $Y_{i}$leftarrow DeconfoundedGeneration($M_{i}$, $X_{i}$, $n$)
     $\{S_{1,1},\dots,S_{N,n}\}$leftarrow EvaluateResponse($L,Y_{i}$)
     $D_{i}$leftarrow ConstructPairs($\{S_{1,1},\dots,S_{N,n}\}$, $X_{i}$, $Y_{i}$)
     $M_{i+1}$leftarrow TrainDPO($M_{i}$, $D_{i}$)
end for
return $M_{k}$
@@ 3.1 Experimental Setup
Evaluation. We evaluate the models from two perspectives, including trustworthiness reflecting the hallucination degree, and helpfulness reflecting the general interaction capability. For trustworthiness, we perform evaluation on five benchmarks:
@@
(3) MHumanEval [58] comprises 146 samples collected from both Object HalBench (50) and MMHal-Bench (96) to provide a more comprehensive evaluation over both long-form description and short-form questions. We only label the overall response-level hallucination rate to control the evaluation cost.
(4) AMBER [51] is a multi-dimensional hallucination benchmark comprising more than 15k samples. We use the discriminative part of AMBER and report the accuracy and F1 metric.
The above trustworthiness evaluations are either limited to common object hallucination, which is mostly eliminated, constrained format (e.g., yes-no choices) or manual labeling. To reliably and automatically assess the trustworthiness of MLLMs under any format, we construct a novel Reliable Free-format Multimodal Benchmark (RefoMB) containing 120 images and 360 instructions covering 8 critical tasks such as mechanical reasoning [35] and image perception [2]. We assess the performance of MLLMs by asking GPT-4 to compare the model response with GPT-4V regarding both trustworthiness and helpfulness. We calculate the trustworthiness win rate and overall win rate based on the GPT-4 evaluation review. Each instruction is paired with a thoroughly written image description as the reference for GPT-4, achieving a notable 96% human agreement. Results on the dev split (99 instructions) are reported in this section to save evaluation costs, we present the test split (261 instructions) results of popular models in the Appendix.
For helpfulness, we adopt two benchmarks: (1) LLaVA Bench [33] is a widely adopted benchmark for assessing multimodal conversation, detailed description and complex reasoning capabilities. We use the in-the-wild version to better capture the capabilities of MLLMs in real-world scenarios. It scores model output against a human annotated image description and reference answer via GPT-4. (2) MMStar [8] is a comprehensive benchmark containing 1500 challenge samples collected from 6 popular multimodal benchmarks [62, 34, 37, 24, 35, 21], covering 6 core capabilities and 18 detailed axes. We report the overall score on this benchmark.
@@
(1) General baselines. We adopt LLaVA 1.5 [31], Qwen-VL-Chat [4], OmniLMM [43], LLaVA-NeXT [32], MiniGemini [28] as representative general baselines. These models are mostly pre-trained on large-scale multimodal data and fine-tuned on high-quality instruction data, achieving strong performance across various multimodal tasks.
(2) Baselines tailored for feedback learning. LLaVA-RLHF [49] trains the reward model on 10k human-labeled preference data and performs proximal policy optimization to train the model on 72k factually augmented data. RLHF-V [58] collects 1.4k fine-grained correctional human feedback and trains the model with the proposed dense direction preference optimization method to reduce hallucination. Silkie [26] utilizes GPT-4V to assess the helpfulness, visual faithfulness and ethical considerations of responses generated by 12 MLLMs on more than 80k multimodal instruction.
(3) Baselines tailored for hallucination reduction without feedback. LURE [70] trains a revisor model to rectify object hallucination in model response. VCD [23] contrasts model logits derived from original and distorted visual input to reduce the over-reliance on statistical bias and unimodal priors. OPERA [18] introduces a penalty term on the model logits. Less-is-more [63] propose a selective end-of-sentence (EOS) special token supervision loss and data filtering strategy.
(4) Proprietary Baseline. We also include GPT-4V [42] as strong reference to evaluate the gap between the open-source models and state-of-the-art proprietary model.
Table 1: Main experimental results. We report hallucination rates in different granularities including response-level (Resp.) and mention-level (Ment.). MHuman.: MHumanEval, Hall.: Hallucination Rate, Trust: trustworthiness win rate, Win.: overall win-rate. The best and second best results are shown in bold and underlined respectively.
Model Size Feedback Object MHuman. MMHal- AMBER LLaVA MMStar RefoMB
HalBench Bench Bench
Resp. $\downarrow$ Ment. $\downarrow$ Resp. $\downarrow$  Score Hall. $\downarrow$ Acc. F1 Overall Overall Trust. Win.
VCD [23] 7B 48.8 24.3 67.1 2.12 54.2 71.8 74.9 65.8 33.8 39.9 16.7
Less-is-more [63] 7B 40.3 17.8 63.7 2.33 50.0 72.4 75.8 60.9 32.9 51.1 16.2
OPERA [18] 7B 45.1 22.3 63.0 2.15 54.2 75.2 78.3 60.3 32.9 33.8 13.1
LURE [70] 7B 27.7 17.3 78.1 1.64 60.4 73.5 77.7 36.9 33.3 29.8 3.0
QWEN-VL [4] 10B 40.4 20.7 61.0 2.76 38.5 81.9 86.4 71.9 34.5 40.9 17.7
LLaVA-NeXT [32] 34B 12.6 6.4 53.4 3.31 34.4 81.4 85.4 77.7 51.6 44.4 35.4
MiniGemini [28] 34B 14.5 8.0 59.6 3.08 38.5 82.6 87.6 79.2 45.5 50.0 36.9
HA-DPO [67] 7B Rule 39.9 19.9 53.4 1.98 60.4 75.2 79.9 67.2 32.9 39.9 17.2
POVID [69] 7B Rule 48.1 24.4 67.8 2.08 56.2 82.9 87.4 62.2 34.3 44.4 13.6
LLaVA-RLHF [49] 13B Human 38.1 18.9 72.6 2.02 62.5 79.7 83.9 61.5 34.2 26.3 17.2
Silkie [26] 10B GPT-4V 27.1 13.4 54.1 3.19 32.3 82.2 87.6 73.2 33.6 38.9 21.2
RLHF-V [58] 13B Human 12.2 7.5 55.5 2.45 51.0 72.6 75.0 51.4 33.2 41.4 17.7
LLaVA 1.5 [31] 7B 53.6 25.2 65.1 2.36 51.0 73.5 77.7 59.7 33.3 30.8 12.1
+ RLAIF-V 7B LLaVA-NeXT 8.5 4.3 37.7 3.06 29.2 76.8 84.5 64.9 35.4 47.5   20.7
OmniLMM [43] 12B 12.2 6.2 50.7 3.14 36.5 86.5 89.5 72.7 39.7 48.0 23.7
+ RLAIF-V 12B OmniLMM 3.3 1.8 29.5 3.36 29.2 88.0 90.9 74.3 40.9 62.1 32.8
GPT-4V [42] - Unknown 13.6 7.3 45.9 3.49 28.1 83.4 87.4 93.1 50.4 50.0 50.0
Implementation Details. We use the Nous-Hermes-2-Yi-34B [3] version of LLaVA-NeXT and the no-RLHF version of OmniLMM [43] as labeler models. For each iteration, we train the model with DPO for 4 epochs, with learning rate 5e-7, beta 0.1, and batch size of 8. We train both RLAIF-V 7B and RLAIF-V 12B for 4 iterations, where we use 4k instructions to collect feedback at each iteration. In summary, it costs 48h and 50h for data collection of 7B and 12B models, and costs 6h and 8h for training separately, using an 8xA100 80G machine.

### v2

# RLAIF-V: Open-Source AI Feedback Leads to Super GPT-4V Trustworthiness

@@ Abstract
Traditional feedback learning for hallucination reduction relies on labor-intensive manual labeling or expensive proprietary models. This leaves the community without foundational knowledge about how to build high-quality feedback with open-source MLLMs. In this work, we introduce RLAIF-V, a novel framework that aligns MLLMs in a fully open-source paradigm. RLAIF-V maximally explores open-source MLLMs from two perspectives, including high-quality feedback data generation for preference learning and self-feedback guidance for inference-time scaling. Extensive experiments on six benchmarks in both automatic and human evaluation show that RLAIF-V substantially enhances the trustworthiness of models at both preference learning and inference time. RLAIF-V 7B reduces object hallucination by 80.7% and overall hallucination by 33.7%. Remarkably, RLAIF-V 12B further reveals the self-alignment potential of open-source MLLMs, where the model can learn from feedback of itself to achieve super GPT-4V trustworthiness.
@@ 1 Introduction
Recent advances in multimodal large language models (MLLMs) mark a significant milestone in AI research [4, 10, 33, 35, 34, 65]. These models are trained on large-scale multimodal corpora and possess profound world knowledge, showing remarkable capabilities in tackling diverse multimodal tasks [27, 39, 43]. However, it has been commonly noticed that MLLMs are prone to confidently generating incorrect content that deviates from human preferences [56, 79, 66, 21]. In order to align MLLMs with human preferences, reinforcement learning from human feedback (RLHF) has been widely used and demonstrates substantial results [56, 66]. However, RLHF depends heavily on labor-intensive human annotations and is consequently hard to cover the widespread misalignment between model and human preferences. Recently, reinforcement learning from AI feedback (RLAIF), which uses the preference collected from labeler models as a proxy of human preference, has shown promising potential as an alternative to RLHF [24].
However, current approaches face two challenges: (1) Infeasible labeler requirement. Existing RLAIF methods, demonstrated at the left top of Figure 1, rely on costly proprietary models to distill feedback from [28, 76, 80]. More critically, this paradigm essentially distills the capability of proprietary models to provide a temporary solution for bridging the performance gap. The community consequently lacks knowledge about how to build high-quality feedback using open-source MLLM labelers of comparable capability, as demonstrated at the left bottom of Figure 1. Simply changing the labeler model from a proprietary model to a weaker open-source model leads to unsatisfactory feedback quality due to their limited capacity [6]. (2) Limited inference-time scaling. Inference-time scaling has drawn great attention from the LLM community and shows promising results [54, 45]. Nevertheless, recent works in MLLMs mainly focus on the preference learning stage to utilize high-quality feedback [66, 78, 74] while omitting the importance of feedback in the inference stage. Besides, aimlessly scaling inference computation budget can hardly contribute to the performance, since accurate feedback guidance plays an important role for effective inference-time scaling.
@@
RLAIF-V addresses these challenges through two key innovations: (1) For high-quality feedback generation, we propose a novel deconfounded candidate response generation strategy for better data efficiency and a divide-and-conquer approach for higher pairwise preference accuracy. The deconfounded strategy accurately exposes the genuine trustworthiness difference within response pairs by generating candidate responses from multiple sampling decoding trials under the same condition. Consequently, confounding factors such as the text style are eliminated, and the feedback focuses on the substantial content of responses. The divide-and-conquer approach decomposes the difficult response-evaluation task into simpler claim-evaluation, which substantially simplifies the task and thus reduces the capacity requirement of labeler models. (2) For inference-time scaling guidance, we propose a novel self-feedback approach based on models aligned with direct preference optimization [48] (DPO). Specifically, we leverage the reward score generated by aligned models as feedback for itself. However, previous works have shown that direct feedback from DPO-aligned models can be biased towards shorter responses due to its objective formulation [49]. We devise a length-normalization strategy to aggregate the token-level scores of each response for bias suppression. Moreover, we also extensively explore the inference-time scaling [54, 45] potential of our RLAIF-V reward on other open-source models, and demonstrate that a single reward model can well generalize to improve the trustworthiness of multiple MLLMs.
@@
Comprehensive experiments on six benchmarks show that RLAIF-V can substantially enhance model trustworthiness without any human or proprietary model intervention. Using feedback from LLaVA-NeXT 34B [35], RLAIF-V 7B significantly reduces the object hallucination on Object HalBench [66] by 80.7%, even surpassing the labeler model. Pushing the limit to an extreme scenario where no stronger models are available, we align OmniLMM 12B [46] with itself as the labeler. Experimental results show that RLAIF-V 12B reduces object hallucination by 76.8% in Object HalBench and overall hallucination by 32.4% in MHumanEval, surpassing GPT-4V by a large margin and revealing the self-alignment potential of open-source MLLMs.
@@
Figure 2: Overview of the RLAIF-V framework. (1) Given the input image and prompt, multiple candidate responses are generated with the deconfounded strategy. (2) Each response is split into atomic claims and assigned with trustworthiness scores separately by an open-source MLLM. (3) During preference learning, the model is aligned under an iterative feedback learning approach which periodically updates the feedback. (4) During inference, performance of the aligned model is further improved by self-feedback guidance.
@@
The contribution of this work can be summarized as fourfold: (1) We present RLAIF-V, a novel framework that aligns MLLMs with open-source feedback. (2) We propose a novel deconfounded and divide-and-conquer approach to generate human-level quality feedback with open-source models. (3) We propose a novel self-feedback guidance for inference-time scaling using aligned MLLMs and devise a simple length-normalization strategy tackling the bias towards shorter responses. (4) We conduct comprehensive experiments to demonstrate the effectiveness of the proposed framework, achieving state-of-the-art performance in trustworthiness among both open-source and proprietary MLLMs. All codes, data, and model weights will be released to facilitate future research.
@@ 2 RLAIF-V
In this section, we first elaborate on how to collect high-quality AI feedback from open-source MLLMs by introducing the response generation and feedback annotation process. Then we introduce the iterative feedback learning stage and the self-feedback guidance for inference-time scaling. An overview of the RLAIF-V framework is demonstrated in Figure 2.
@@ 2.2 Feedback Annotation
Combine. After obtaining the quality assessment of each claim, we finally combine them into the score of the whole response. For each response, we denote the number of claims having $p_{no}>p_{yes}$ as $n_{rej}$, measuring how many incorrect claims are recognized by the labeler model. We use $-n_{rej}$ as the final score $S$ of the response, where a higher score indicates less incorrectness of the content. Given the score of each response, we can now construct a preference dataset for training. For each instruction $x$, we keep all response pairs $(y,y^{\prime})$ such that $S>S^{\prime}$ and choose the higher score response $y$ as the preferred response. To save the training cost, we randomly sample at most 2 pairs for each instruction and we find such a filtering process only causes minor performance drops. To prevent the potential length bias, we drop pairs in which $y_{w}$ is too short before training to ensure the average length difference of $y_{w}$ and $y_{l}$ is less than one word.
2.3 Iterative Feedback Learning
@@
DPO is widely used to align MLLMs with human preferences. However, naive DPO faces the distribution shift problem, i.e., the preference data is static during the training process while model output distribution is constantly shifting [14]. As a result, the data distribution might deviate from the expected feedback distribution and cause suboptimal alignment results.
@@
We follow [17] to train the model iteratively. Specifically, we select $N$ multimodal instructions at the beginning of each iteration and leverage the deconfounded strategy to generate $n$ candidate responses for each instruction with the latest instruction model $M_{i}$. We assign each response with a trustworthiness score through the divide-and-conquer approach using the labeler model $L$ and construct comparison pairs $D_{i}$ for training. Then we train the $M_{i}$ with direct preference optimization on $D_{i}$ to get $M_{i+1}$, which is used as the instruction model of the next iteration. In this way, the feedback distribution can be updated in an iterative manner, resulting in better learning efficiency.
2.4 Self-Feedback for Inference-time Scaling
After iterative learning on diverse high-quality feedback, the MLLM itself is not only a trustworthy policy model but also a reward function by the optimization objective of DPO [48] and the reward is formulated as:
$\begin{split}r(y)&=\beta\log{\frac{\pi_{\theta}(y)}{\pi_{\text{ref}}(y)}}=\beta\sum_{t}^{T}{\log{\frac{\pi_{\theta}(y_{t}|y_{<t})}{\pi_{\text{ref}}(y_{t}|y_{<t})}}},\end{split}$ (1)
where $\beta$ is a parameter controlling the deviation from the base reference policy $\pi_{ref}$, $y$ is the response token sequence, $T$ is the length or response and $\pi_{\theta}$ is the model after DPO training. We hide the prompt condition $x$ in equations for simplicity. Previous works have shown that DPO-aligned reward $r(y)$ can be biased towards shorter responses due to its objective formulation [49]. We tackle this bias by averaging all token-level scores to get the final response score $r(y)=\frac{\beta}{T}\log{\frac{\pi_{\theta}(y)}{\pi_{\text{ref}}(y)}}$.
We then use the normalized reward as self-feedback guidance for inference-time scaling. Specifically, we follow [54] to perform best-of-N (BoN) selection based on multiple sampled responses of the same prompt. Specifically, we choose the response with the highest score among the N candidate responses as the model prediction. To amplify the candidate response diversity, we follow existing works [54] to apply commonly used nucleus sampling [19] for decoding.
@@ 3 Experiments
Model Size Feedback Object MHum. MMHal- AMBER MM- RefoMB
HalBench Bench Star
Rsp. $\downarrow$ Men. $\downarrow$ Rsp. $\downarrow$ Score Hall. $\downarrow$ Acc. F1 Avg. Trust. Win.
VCD [25]
(CVPR’24)
7B 48.8 24.3 67.1 2.12 54.2 71.8 74.9 33.8 39.9 16.7
Less-is-more [71]
(ACL’24)
7B 40.3 17.8 63.7 2.33 50.0 72.4 75.8 32.9 51.1 16.2
OPERA [21]
(CVPR’24)
7B 45.1 22.3 63.0 2.15 54.2 75.2 78.3 32.9 33.8 13.1
CCA-LLaVA [64]
(NeurIPS’24)
7B 46.7 23.8 68.5 1.92 61.5 77.7 81.9 32.1 41.9 21.7
Qwen-VL-Chat [4]
(arXiv’23)
10B 40.4 20.7 61.0 2.76 38.5 81.9 86.4 34.5 40.9 17.7
LLaVA-NeXT [35]
(arXiv’24)
34B 12.6 6.4 53.4 3.31 34.4 81.4 85.4 51.6 44.4 35.4
MiniGemini [31]
(arXiv’24)
34B 14.5 8.0 59.6 3.08 38.5 82.6 87.6 45.5 50.0 36.9
HA-DPO [76]
(arXiv’23)
7B Rule 39.9 19.9 53.4 1.98 60.4 75.2 79.9 32.9 39.9 17.2
POVID [78]
(arXiv’24)
7B Rule 48.1 24.4 67.8 2.08 56.2 82.9 87.4 34.3 44.4 13.6
LLaVA-RLHF [56]
(arXiv’23)
13B Human 38.1 18.9 72.6 2.02 62.5 79.7 83.9 34.2 26.3 17.2
Silkie [28]
(EMNLP’24)
10B GPT-4V 27.1 13.4 54.1 3.19 32.3 82.2 87.6 33.6 38.9 21.2
RLHF-V [66]
(CVPR’24)
13B Human 12.2 7.5 55.5 2.45 51.0 72.6 75.0 33.2 41.4 17.7
AMP-MEG [74]
(NeurIPS’24)
13B Rule 31.7 20.6 54.8 3.08 36.5 79.5 84.6 34.8 30.3 14.6
LLaVA 1.5 7B 54.5 27.8 67.1 1.86 63.5 73.5 77.7 33.3 36.9 16.2
+ RLAIF-V 7B LLaVA-NeXT 10.5
+44.0
5.2
+22.6
44.5
+20.6
2.95
+1.1
32.3
+31.2
76.8
+3.3
84.5
+6.8
35.4
+2.1
47.2
+10.3
22.5
+6.3
+ RLAIF-V BoN 7B  LLaVA-NeXT 6.8
+3.7
3.8
+1.4
39.7
+4.8
3.07
+0.1
28.1
+4.2
N/A N/A N/A 55.7
+8.5
24.4
+1.9
OmniLMM 12B 19.4 10.9 52.7 3.06 36.5 86.5 89.5 39.7 44.7 18.5
+ RLAIF-V 12B self 4.5
+14.9
2.9
+8.0
35.6
+17.1
3.15
+0.1
32.3
+4.2
88.0 +1.5 90.9 +1.4 40.9
+1.2
58.1
+13.4
28.3
+9.8
+ RLAIF-V BoN 12B self 4.5 +0.0 2.6 +0.3 29.5
+6.1
3.44
+0.3
26.0 +6.3 N/A N/A N/A 62.9 +4.8 30.3
+2.0
GPT-4V [43] - Unknown 13.6 7.3 45.9 3.49 28.1 83.4 87.4 50.4 50.0 50.0
Table 1: Main experimental results. We report hallucination rates in different granularities including response-level (Rsp.) and mention-level (Men.). MHum.: MHumanEval, Hall.: Hallucination Rate, Trust.: trustworthiness win rate, Win.: overall win-rate. The best results are shown in bold. BoN: apply RLAIF-V 7B and RLAIF-V 12B self-feedback for best-of-N, we sample 32 and 16 samples respectively to control evaluation cost. N/A: multi-choice and yes-no question do not have BoN results since these questions only requires single token.
@@ 3.1 Experimental Setup
Data ObjHal. AMBER
Rsp. $\downarrow$ Men. $\downarrow$ Acc. F1
RLHF-V [66] 28.5 12.3 76.4 84.6
RLAIF-V 10.1 4.7 80.1 86.1
w/o deconfounding 25.7 11.8 73.3 83.0
Evaluation. We evaluate models from two perspectives, including trustworthiness reflecting the hallucination degree, and helpfulness reflecting the general capability. For trustworthiness, we perform evaluation on five benchmarks:
@@
(3) MHumanEval [66] comprises 146 samples collected from both Object HalBench (50) and MMHal-Bench (96) to provide a more comprehensive evaluation over both long-form description and short-form questions. We only label the response-level hallucination rate to control the cost.
(4) AMBER [58] is a multi-dimensional hallucination benchmark comprising more than 15k samples. We use the discriminative part and report the accuracy and F1 metric.
The above trustworthiness evaluations are either limited to common object hallucination, which is mostly eliminated, constrained format (e.g., yes-no choices) or manual labeling. To reliably and automatically assess the trustworthiness of MLLMs under any format, we construct a novel Reliable Free-format Multimodal Benchmark (RefoMB) containing 120 images and 360 instructions covering 8 critical tasks such as mechanical reasoning [38] and image perception [2]. Following [34], we assess the performance of MLLMs by comparing the model response with GPT-4V response regarding both trustworthiness and helpfulness. We calculate the trustworthiness win rate and overall win rate based on the evaluation review. Each instruction is paired with a thoroughly written image description as the reference, achieving a notable 96% human agreement. Results on the dev split (99 instructions) are reported in this section to save evaluation costs, we present more details and the test split (261 instructions) results in the Appendix.
Data Agree. ObjHal. AMBER
Rsp. $\downarrow$ Men. $\downarrow$ Acc. F1
VL-Feedback [28] 92.3% 37.9 21.0 72.8 82.6
RLAIF-V 96.7% 20.6 10.4 80.5 86.0
     w/o d&c 66.7% 53.1 26.2 73.5 77.7
     w/ smaller labeler 90.0% 33.0 17.5 75.2 78.2
Table 3: Performance comparison of different feedback collection methods. We conduct the experiment on various labeler models. ObjHal.: Object HalBench. smaller labeler: OmniLMM 12B, Agree.: Human agreement of the constructed pairs, d&c: divide-and-conquer strategy. VL-Feedback collects high-quality feedback from GPT-4V.
For helpfulness, we adopt MMStar [8], which is a comprehensive benchmark containing 1500 challenge samples collected from 6 popular multimodal benchmarks [70, 36, 37, 26, 38, 23], covering 6 core capabilities and 18 detailed axes. We report the overall score on this benchmark.
@@
(1) General baselines. We adopt LLaVA 1.5 [33], Qwen-VL-Chat [4], OmniLMM [46], LLaVA-NeXT [35], MiniGemini [31] as representative general baselines.
(2) Baselines tailored for feedback learning. RLHF-V [66] collects fine-grained correctional human feedback and trains the model with dense direction preference optimization. Silkie [28] utilizes GPT-4V to collect feedback. POVID [78] and AMP-MEG [74] apply heuristic rules to pair responses generated under difference condition.
(3) Baselines tailored for hallucination reduction without feedback. VCD [25] contrasts model logits derived from original and distorted visual input to reduce the over-reliance on statistical bias and unimodal priors. OPERA [21] introduces a penalty term on the model logits. Less-is-more [71] proposes a selective end-of-sentence (EOS) special token supervision loss and data filtering strategy. CCA-LLaVA [64] mitigates hallucination by applying a novel concentric causal attention.
(4) Proprietary baseline. We also include GPT-4V [43] as strong reference to evaluate the gap between the open-source models and proprietary models.
Implementation Details. We use the Nous-Hermes-2-Yi-34B [3] version of LLaVA-NeXT and the no-RLHF version of OmniLMM [46] as labeler models. For each iteration, we train the model with DPO for 4 epochs, with a learning rate 5e-7, beta 0.1, and batch size of 8. We train both RLAIF-V 7B and RLAIF-V 12B for 4 iterations, where we use 4k instructions to collect feedback at each iteration. In summary, it costs 48h and 50h for data collection of 7B and 12B models, and costs 6h and 8h for training separately, using an 8xA100 80G machine. For the best-of-N setting, we sample 32 and 16 candidate responses for RLAIF-V 7B and RLAIF-V 12B respectively to control the evaluation cost.